OpenStudy (anonymous):

re write in vertex form Y=x^2-6x+4

OpenStudy (anonymous):

Y=x^2-2(3)x +9 +4 -9 y+9=(x-3)^2 vertex 3,-9

OpenStudy (anonymous):

completely confused

OpenStudy (tkhunny):

Completely the Squarely! Think on this idea: \((a+b)^{2} = a^{2} + 2ab + b^{2}\). It's particularly the b-part that should sink into your brain.

OpenStudy (anonymous):

listen im sorry but i have a hard time in algebra 2 im pretty sure this explination is good but i dont understand any of this because im homeschooled and i have a hard time in math and im failing even though i have straight a's in every other class :(

OpenStudy (tkhunny):

?? You're supposed to think about it, not just suddenly manage to get it. The study of mathematics should involve a thought process, not just a memorization process. We have this: \(x^{2} - 6x + 4\) How does that relate to: \(a^{2} + 2ab + b^{2}\) a in the second expression takes the place of xe in th first expression. Then, how does this: \(x^{2} - 6x + 4\) Relate to this? \(x^{2} + 2xb + b^{2}\) Seriously, stare at those b-parts... If "2xb" relates to "-6x", then 2b = -6 and b = -3. Now what? We are learning to "Complete the Square"! What could be more fun?!

OpenStudy (anonymous):

this kinda makes sense lbut i still cant figure it out

OpenStudy (tkhunny):

You have '\(b\)' and you need '\(b^{2}\)'. How do you propose we should find \(b^{2}\)?

OpenStudy (anonymous):

b*b

OpenStudy (anonymous):

what i dont understand is the 2ab and the 6x